Minutes of the Trigger Operation Meeting (May 9, 2006)
Attendees: Rainer Bartoldus, Al Eisner, Fang Fang, Selina Li, Rahmat Rahmat, Dong Su, Dan Walker
- PEP has been having some troubles with BPM's and their magnets, so we've only been taking some data with low currents and with access for a few times for the past week. No operational problems with the trigger.
- Trigger experts should remember to renew/sign their Trigger JHAM. Note that system worker JHAMs expire after a year.
- TRG status report will be given tomorrow at IR2 after the detector operation meeting. A link to the talk is provided here. If there is anything you would like to be included in the report, please send Selina an email.
- Walt Innes mentioned that the old DCT crates should be disassembled before the beginning of July. The earlier we can have the racks free up the better.
- No news from Marc in regarding to progress on the firmware issue yet. We are currently using the test_9 version of the firmware and still have some small level of TSF damages. This should NOT be a long term solution, and we are hoping to get another version from Marc soon.
Data Quality and Monitoring
- Rahmat is still working on getting the uncertainties on the lumi ratio right. See the HN posted and its follow-ups.
Trickle Injection Monitoring
- We need to work on providing a stand-alone monitoring check for Trickle Injection for the validation review. Binning and projection for slices of the 2-D histograms are briefly discussed. We expect to have more and better discussions of this topic next week.
This page is created by Selina Li (firstname.lastname@example.org); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.